A*B is defined to be 10A-B squared. Evaluate (2*(-1))*5
Answer by stanbon(57967)

A*B is defined to be 10A-B squared. Evaluate (2*(-1))*5
------------
2*(-1) = 10(2)--1 = 21
---
21*5 = 10*21-5 = 210-5 = 205
==============================
Cheers,
Stan H.